OpenStudy (anonymous):

the total hang time T, in seconds, is given y T(h)= √h/2 where h is the mximun height if a person jumps 5 feet off the ground, determine fr how mny seconds the person is in the air

OpenStudy (anonymous):

is it \[\frac{\sqrt{h}}{2}\] or \[\sqrt{\frac{h}{2}}\]